Homework problem- this has to be solved using systems of equations, not a graph. I have to show all work, so I'll definitely need the equations used:
HIS taxi company charges 1 dollar per mile, with a $3 service fee. HERS taxi company charges 75 cents per mile, with a $5 service fee. At what distances do the two taxi companies charge the same amount?
Thanks!!

Answer by stanbon(75887)

HIS taxi company charges 1 dollar per mile, with a $3 service fee. HERS taxi company charges 75 cents per mile, with a $5 service fee. At what distances do the two taxi companies charge the same amount?
----
HIS Equation:
A(x) = 3 + 1x where "x" is the miles driven
====
HERS Equation:
B(x) = 5 + 0.75x where "x" is the miles driven
-----
Solve A(x)=B(x):
---
3+x = 5+0.75x
0.25x = 2
x = 8 miles
-------------------
Cheers,
Stan H.
-------------